OpenStudy (anonymous):

Combine like terms to simplify the expression: 7x+3y-2+6x+1+y^2 A: y^2+3y+13x-1 B: 3y^3+13x^2+1 C: 2y^3+13x-1 D: y^2+3y+13x^2+1

OpenStudy (mathstudent55):

Do you know what like terms are?

OpenStudy (anonymous):

yes @mathstudent55

OpenStudy (anonymous):

just plz tell me the answer @mathstudent55

OpenStudy (mathstudent55):

What is 7x + 6x?

OpenStudy (mathstudent55):

What is -2 + 1?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

I KNOW its A or D!! @mathmath333

OpenStudy (anonymous):

Please help me :( @swagmaster47

OpenStudy (anonymous):

He just answered your question! There is only one option with -1 as an option :)

OpenStudy (mathstudent55):

\(7x+3y-2+6x+1+y^2\) \(=y^2 + 3y \color{red}{+7x + 6x} \color{green}{-2 + 1} \) \(=y^2 + 3y \color{red}{+13x} \color{green}{-1} \)